OpenStudy (anonymous):

find a degree 3 polynomial with zeros -2, 1, 5 and go through point (0, -3)

OpenStudy (saifoo.khan):

Did u tried that?

OpenStudy (anonymous):

yeah help

OpenStudy (anonymous):

x^3-4x^2-7x+10

OpenStudy (anonymous):

how did you figure that out

OpenStudy (anonymous):

Well, since -2,1,5 are zeros. Then they must be... x+2=0 x-1=0 x-5=0 Since all those would equal to 0. Then you just multiply the three. (x+2)(x-1)(x-5). And you get the answer.